OpenStudy (anonymous):

An early method of measuring the speed of light makes use of a rotating slotted wheel. A beam of light passes through one of the slots at the outside edge of the wheel, travels to a distant mirror, and returns to the wheel just in time to pass through the next slot in the wheel. One such slotted wheel has a radius of 8.0 cm and 950 slots at its edge. Measurements taken when the mirror was l = 500 m from the wheel indicated a speed of light of 3.0 105 km/s.(a) What was the (constant) angular speed of the wheel(rad/s)? (b) What was the linear speed of a point on the edge of the wheel (m/s)

OpenStudy (anonymous):

This is my work so far (a) d=2*L d=2*500m d=1000m t=L/c t=500m/(3*10^8m/s) t=0.000002s angular speed=rw Where do I go from here?